Circular No. 7799

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 A3 (LINEAR) An object reported as asteroidal by LINEAR (discovery observation

below) has been noted by several CCD observers to be cometary: Jan. 19.9 UT, slightly diffuse, m_{1} = 17.3 (J. Nomen, Barcelona, Spain); 20.6, 10′′ coma, m_{1} = 17.3 (J. Broughton, Reedy Creek, Qld.); 21.2, fuzzy coma of diameter ∼ 10′′ (R. Dyvig, Quinn, SD); 21.2, m_{1} = 17.0, 22′′ tail in p.a. 250^{o} (R. Fredrick, R. Trentman, and R. Gruenke, Louisburg, KS).

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 13.22750  6^{h}42^{m}26.^{s}93  −13^{o}34′44′′.8  17.6

MPEC 2002-B18 has astrometry, orbital elements (below) and ephemeris:

T = 2002 May 3.404 TT  ω = 330.^{o}334
Ω = 137.279 }2000.0
q = 5.08347 AU  i = 47.389

SUPERNOVA 2002I IN IC 4229 D. Weisz and W. D. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7514) discovery of another apparent supernova (mag ∼ 17.3) on unfiltered KAIT images taken on Jan. 19.5 and 20.5 UT. SN 2002I is located at α = 13^{h}22^{m}26^{s}.84, δ = −2^{o}25′04′′.1 (equinox 2000.0), which is 11′′.1 east and 2′′.6 north of the nucleus of IC 4229. A KAIT image taken on 2001 June 16.3 showed nothing at this position (limiting mag ∼ 19.0)

IM NORMAE H. W. Duerbeck, Vrije Universiteit (VU), Brussels; R. Baptista, Depar-

tamento de Física, Universidade Federal de Santa Catarina, Florianopolis; C. M. Dutra, Departamento de Astronomia, Universidade de São Paulo; and C. Sterken, VU, report low- and medium-resolution spectroscopy of this recurrent nova on Jan. 14.4, 15.4, and 17.4 UT with the European Southern Observatory 1.52-m telescope at La Silla: “The lines Hβ–Hϵ show P-Cyg profiles, with an average FWHM of 1150 km/s and a radial velocity of +400 km/s for the emission component, and a radial velocity of −860 km/s for the weak absorption component. The spectrum is dominated by emission of Fe II (multiplets 27, 38, 42, 49) in the blue and of O I (1, 4) and Ca II (2) in the red. This is the first recurrent nova belonging to the Fe II class (see Williams 1992, A.J. 104, 725).”

2002 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green